# Forming the local starburst galaxy Haro 11 through hydrodynamical merger simulations

### v1

@@ 3.3 Strategy for tweaking initial conditions
Other than the velocity, the mass ratio between the galaxies has a large impact on the orbit. However, compared to the velocity the masses are not as easy to fine-tune and has additional effects on other observable. For example, the SFR depends on the orbit but the amount of stars formed in the individual galaxies depends on their halo properties in a non-linear way. Thus, the mass ratio was primarily changed in early tests.
@@
Figure 2: The column density of gas (top row) and stars (bottom row) for the large-scale structure of the galaxy are shown at various time outputs, spanning from the initial interaction to the point where the galaxy most closely resembles Haro 11. The plots marked with Roman numbering correspond to the times in the SFR plot in Figure 7.
@@ 3.4 Analysing the simulation output
The total runtime of the fiducial simulation is 1280 Myr, which is just after the centres of the two galaxies fully coalesce. We analyse the galaxy at the time output we deem to be closest in terms of SFR and morphology to the Haro 11 galaxy. This, current observed, time is set as 0 Myr and is as the galaxies are undergoing their second interaction. The morphology and kinematics we calculate are largely invariant on the choice of time by several Myr, but the SFR can vary by a factor of 2-4 within only a few Myr. After $\gtrsim 10$ Myr, the morphology (and SFR) is notable different to Haro 11.
The exact morphology and kinematics of the inner galaxy is dependent on the viewing angle we adopt. Our approach is to find an angle by eye that improves the match for the inner morphology (i.e. relative position between knots and the ’ear’) while keeping the tidal tail at roughly the correct relative angle to the inner morphology. The fiducial viewing angle is in-between the $z$ and $y$ axis of the simulation (specifically, $\hat{n}=[0.26,-0.64,0.72]$). The morphology is insensitive to a change in angles within a subtended angle $\sim 10^{\circ}\times 10^{\circ}$, with some directions being less or more affected.
@@
[…]
Figure 4: The stellar tidal tail from observations (left) and simulations (right). The simulation has green contours showing the gas distribution and the red contour highlights the inner region. The observations are in the B-band with VLT/FORS (see details in Section 4.1), with green contours showing the stellar distribution.
Figure 5: Projections maps of the inner part of the galaxy at the output where the galaxy most closely resembles Haro 11.
Figure 6: Projection map of the stellar mass weighted by their inverse stellar age, $M_{*,}=M_{*}\times(t/4\,{\rm Myr})^{-1.25}$, to represent young, bright stars. The red/green contours correspond to the projected stellar/gas column density. Combined with Figure 1, in which knot A is not readily observed due to this extinction, this map demonstrates that bright stars form in knot A but are obscured by dust.
@@ 4 The fiducial Haro 11 model
Table 3: Properties of the galaxy from the fiducial Haro 11 simulation, analysed at current time, when the morphology and SFR best match observations, for a volume of 60 x 60 x 60 kpc3 around the central part of the galaxy. We separate between ionised gas around the galaxy ($<10$ kcc) and ionised CGM gas, as the latter has no observational comparison. The CGM is, thus, excluded from the gas fraction and total gas mass.
Stellar mass [$10^{9}{\rm\,M_{\odot}}$] 16.49
Total gas mass [$10^{9}{\rm\,M_{\odot}}$] 4.33
@@
Molecular gas mass [$10^{9}{\rm\,M_{\odot}}$] 1.35
Neutral gas mass [$10^{9}{\rm\,M_{\odot}}$] 1.06
Ionised gas mass [$10^{9}{\rm\,M_{\odot}}$] 1.77
Ionised CGM gas mass [$10^{9}{\rm\,M_{\odot}}$] 1.90
SFR [${\rm\,M_{\odot}\,yr^{-1}}$] 43
@@ 4.1 Tidal tail
The first step of analysing the simulation is to visualise the merger process and the tidal tail, as it is one of the strongest signs of Haro 11 being a merger. In Figure 2 we present several projection maps of the gas and stars as part of a timeline of the simulation. These maps follow the formation of the tidal tail, from the the first close interaction until they reach the current observed time of Haro 11, which is $\sim 10$ Myr into the 2nd interaction; a time period of roughly $\sim 200$ Myr. Furthermore, the maps have Roman numbers annotated, corresponding to time markers in the figure of SFR as a function of time in Section 4.3, which highlights the burst of star formation during the close interactions.
@@
We next present the detection of a stellar tidal tail in Haro 11, the first to confirm the existence of a stellar component to the tidal tail. The observation was made in the B-band with VLT/FORS for a net integration time of 6240 s and reduced in a standard manner [^2] Applying bias subtraction, cosmic ray removal, flat fielding and atmospheric extinction correction.. The image is presented on the left plot in Figure 4, along with an image of the stars in our simulations on the right. Furthermore, there is a much shorter, parallel component, to the tail and a short feature extending to the upper right (NW). The latter two features are contaminated by the extended PSFs from bright stars, but are real. In addition, an archival HST/WFC3/IR/F160W image confirms the existence of these features at low S/N but the field of view is too small to fully map the Haro11 tidal system. The morphology of the tidal tail agrees with our simulation, but also the small feature extending in the upper right from Haro 11 can be found in our fiducial galaxy.
The stellar tidal tail is slightly shorter than the HI component ($\sim$30 kpc compared to $\sim$40 kpc). This is likely due to a difference in scale radius or truncation between the two components, i.e. the gas disc extending further. We briefly explored the effects of truncation on the disc in Section 5.1. Additionally, the image of the stellar tidal tail shows clear shell structures around the galaxy. The simulation shows signs of these shells as well, which is encouraging for our model but is not useful as a constraint.
@@ 4.2 Inner morphology
In order to put our simulations into context, we first present an observations of the Haro 11 galaxy on the left in Figure 1. This was presented in Adamo et al. (2010) using HST in three wavebands filters; in particular they use 220W and 814W with the ACS/HRC instrument, and 435W with ACS/WFC. Additionally, the right plot of the figure shows a mock observation of our simulation using the post-processing radiative transfer code RASCAS (Michel-Dansac et al., 2020) to obtain the stellar continuum and apply the same HST filter profiles as the observations. The key morphological features of Haro 11 are annotated: three distinct stellar knots (A, B, C) and a dusty arm going upwards from knot A to the right of knot B (the ’ear’). All of these morphological features can be identified in the mock observation as well, but knot A is not as readily apparent, which is discussed below.
The inner gas and stellar morphology, as well as temperature, of the galaxy at the current observed time is presented as projection maps in Figure 5. The high density gas around knot B and A correspond to areas of high stellar density. The disc arm with young stars is seen in the right-side of the plots is reminiscent of the ear observed in Haro 11. Compared to observations, the positions of these features are slightly mismatched in position relative to each other, but are broadly in the expected positions. Additionally, the dense gas that is overlapping with the stellar core on the left (knot C) is behind the core.
The main reason that knot A is not observed in our mock observations is due to the dusty arm covering the stars, but we can infer its presence from the stellar mass and young population. In Figure 6 we show a stellar map weighted by the inverse stellar age, as $M_{*,}=M_{*}\times(t/4\,{\rm Myr})^{-1.25}$, which is an approximate bolometric luminosity weighting; this highlights the presence of young, luminous stars. This figure shows that there is significant mass of young stars around knot A, which would be visible in the absence of extinction. Additionally, in Figure 8 we present a projection map of the median stellar ages, from which it is also evident that all of the knots, including A, contain many young stars. From this, we surmise that knot A likely reside within the disc arm (which also makes up the ear) of the prograde progenitor galaxy, which has its centre at knot B.
@@ 4.3 Star formation rate
Observations show that Haro 11 is currently experiencing a burst of star formation, which is attributed to an undergoing merger. We present the SFR as a function of time in Figure 7, annotated with Roman numbers as timestamps to the timeline images in Figure 2 and with shaded areas for the time of pericentre passages. The figure shows two peaks in the SFR that are associated with close interaction between the two galaxies. The SFR burst is thus likely caused by interaction effects, e.g. gas compression due to tidal forces, cloud-on-cloud collision, tidal torques that drive gas towards the centres of the galaxies (e.g. Renaud et al., 2014b, 2019).
The current day SFR from observations match with the observed values of our fiducial simulation. Furthermore, the starburst periods in our simulations can be readily compared to observations of stellar populations with distinct ages, which helps constrain the star formation history. The peaks indicate that these stellar populations should have ages of roughly $200-250$ Myr, from the first interaction, and a few Myr, from the current interaction. We will explore this further in the next section by comparing stellar analysis models of the populations in the three knots.
Figure 7: The star formation rate of the simulation. The time axis is set that the current observed time is at 0 Myr. Only the last $\sim 500$ Myr are shown, as there is no interaction between galaxies before this and the galaxies have reached a roughly constant SFR (see Section 3.4 for details). Specific time periods are marked with Roman numbers, corresponding to the projection maps with same number in Figure 2. The grey shaded areas are the estimate pericentre periods of interaction. The starbursts at 0 Myr and 250 Myr correspond to close, gravitational interactions. The diamond marks the time at which the galaxy best resembles Haro 11, following the criteria in Section 3.2, and the square represent the SFR range mentioned in Section 2. The galaxy is in the middle of its second starburst event, with a ${\rm SFR}=43{\rm\,M_{\odot}\,yr^{-1}}$.
@@ 4.4 Knot properties
The distribution of mass and median ages of the clusters, divided into their knot affiliation, are presented in Figure 9. As mentioned in Section 4.3, we notice two massive populations of star clusters that form during the starburst peaks that coincide with the close interactions. These groups of population ages, $\sim 10^{6}$ Myr and $\sim 10^{8}$ Myr, have been found in stellar population synthesis models of Haro 11 (Chandar et al., 2023; Papaderos & Östlin, 2023). We recover in the simulation star clusters as massive as those found in the observations at the youngest age, while the majority of the old population would remain undetected at masses below a few times $10^{5}{\rm\,M_{\odot}}$; as can be seen from the 90% completeness limit from Sirressi et al. (2022), presented as a dark line. The most massive clusters found in our simulations $\sim 10^{8}{\rm\,M_{\odot}}$ are not found in the observations, but could stem from insufficient mass resolution, the cluster identification, or other, more spurious, effects. We plan to, in Ejdetjärn et al. (in prep), perform more extensive analysis of the cluster formation within our high-resolution simulations of the same setup.
@@ 4.5 Kinematics
In Ö15, the authors made a comparison between the position-velocity (PV) diagram of the Antennae galaxies and Haro 11, which they noted had a similar shape and velocity magnitude. They acquired data using the CIGALE instrument and archived data for the Antennae. Their data covered a straight line going from knot C to A, and extending a few kpc beyond them. We follow their method and present our own PV diagram in Figure 11. The $x$-axis is determined as the projected distance along the dark line on the left plot in Figure 10. The left/right plot in the figure shows the simulation/observational data, which have very similar shapes with peaks and valleys at roughly the same distance intervals. We highlight these similarities, with a broad grey line in both plots, which is the mean velocity of the simulation at each distance in bins of 0.5 kpc. This figure also highlights that several of the simulations cells attain much higher speeds, as much as twice as fast, as the observations. However, the mean velocity (grey line) follows the same sine-like shape and are within the limits of the observational data.
@@ 5.1 Producing one tidal tail
Additionally, truncating the gas disc also has an effect on the resulting morphology of the tidal tails. In the centre row of Figure 12 we show an initial galaxy with retrograde and prograde motions (left), a thicker disc for the retrograde (middle), and a truncated disc for the retrograde (right). The thicker disc can be seen to yield a slightly weaker 2nd tail and have more tidal features surrounding it. The truncated disc also shows a more diffuse 2nd tail and possibly some small changes to the main tail.
From our simulation tests, we learned that the length of the tidal tail puts a constraint on the time since the first interaction, when the tail first forms. In order for the gas tail to reach the observed 40 kpc, it would need roughly 200 Myr to expand. For comparison, a shorter interaction time of $\sim 100$ Myr would yield a length of $\sim$20 kpc. This time might be a lower limit, as the observed expansion speed of tidal tail is slightly lower in the HI observations. Furthermore, the presence of a tidal tail this long also substantiates that Haro 11 is at least in its 2nd close interaction.

### v2

@@ 3.3 Strategy for tweaking initial conditions
Figure 2: The column density of gas (top row) and stars (bottom row) for the large-scale structure of the galaxy are shown at various time outputs, spanning from the initial interaction to the point where the galaxy most closely resembles Haro 11. The plots marked with Roman numerals correspond to the times in the SFR plot in Figure 7.
Other than the velocity, the mass ratio between the galaxies has a direct impact on the orbit. However, compared to the velocity, the masses are not as easy to fine-tune and have additional effects on other observables. For example, the SFR depends on the orbit but the amount of stars formed in the individual galaxies depends on their halo properties in a non-linear way. Thus, the mass ratio was primarily changed in early tests.
@@
Our strategy at each iteration was as follows. Most tweaks were done one feature at a time. We first improved upon the match with the large-scale and general properties, i.e. tidal tail properties (morphology, kinematics, stripped mass), stellar populations, and SFR, before focusing on the inner region. This was done by tweaking the initial velocities, inclination, and relative mass ratio. For the inner part of the merger, we fine-tuned the stellar and gaseous morphology of the knots and the "ear" by slightly shifting the velocity, inclination, and galactic properties; this of course has a small impact on the tidal tail properties, but these were minor relative to the improved match in the inner region.
As observations of the tidal tail position already aligned well with the $z$-axis of the fiducial setup, the viewing angle only required a relatively small adjustment from this axis, and was done last. The choice of viewing angle was done to position the knots at the expected positions in the projected 2D space (relative to each other and the tidal tail), while also keeping the tail position at a good match with observations.
@@ 3.4 Analysing the simulation output
The total runtime of the fiducial simulation is 1280 Myr, which is just after the centres of the two galaxies fully coalesce. We analyse the galaxy at the time output we deem to be closest in terms of SFR and morphology to the Haro 11 galaxy. We define the snapshot that best matches the observations (i.e. current day) as 0 Myr, during the second close interaction. The morphology and kinematics we calculate are largely invariant on the choice of time by several Myr, but the SFR can vary by a factor of 2-4 within only a few Myr. After $\gtrsim 10$ Myr, the morphology and SFR are notably different to Haro 11.
The exact morphology and kinematics of the inner galaxy is dependent on the viewing angle we adopt. Our approach is to find an angle that improves the match for the inner morphology (i.e. relative position between knots and the ’ear’) while keeping the tidal tail at roughly the correct relative angle to the inner morphology. The fiducial viewing angle is in-between the $z$ and $y$ axis of the simulation (specifically, $\hat{n}=[0.26,-0.64,0.72]$). The morphology is insensitive to a change in angles within a subtended angle $\sim 10^{\circ}\times 10^{\circ}$, with some directions being less or more affected.
@@
Figure 3: Velocity maps of the HI tidal tail. Left panel shows the 21-cm line, in emission only, of Haro 11 with MeerKAT (Le Reste et al., 2024, see their paper for specifics of the observation and data reduction). Right panel shows the simulation; the green contour outlines the central stellar component, and the black contour outlines the gas column density exceeds 10% of its maximum value.
Figure 4: The stellar tidal tail from observations (left) and simulations (right). The observations are in the B-band with VLT/FORS (see details in Section 4.1), with cyan contours showing the stellar distribution for different thresholds. The simulation has green contours that highlight the gas distribution of the tidal tail and the inner region.
Figure 5: Projections maps of the inner part of the galaxy at the output where the galaxy most closely resembles Haro 11. The column densities are calculated as the sum of mass within a pixel and then divided by the pixel area. The temperature is the mass-weighted mean. Green/red contours correspond to the most dense regions of gas/stars.
Figure 6: Projection map of the stellar mass-weighted by their inverse stellar age, $M_{\rm*,\,young}=M_{*}\times(t/4\,{\rm Myr})^{-1.25}$, to represent young, bright stars. The red/green contours correspond to the projected stellar/gas column density. Combined with Figure 1, in which knot A is not readily observed due to this extinction, this map demonstrates that bright stars form in knot A but are obscured by dust.
@@ 4 The fiducial Haro 11 model
Table 3: Properties of the galaxy from the fiducial Haro 11 simulation, analysed at current time, when the morphology and SFR best match observations, for a volume of 60 x 60 x 60 kpc3 around the central part of the galaxy. We separate between ionised gas around the galaxy (<10<10 kpc) and ionised CGM gas, as the latter has no observational comparison. The CGM is, thus, excluded from the gas fraction and total gas mass.
Stellar mass $10^{9}{\rm\,M_{\odot}}$] 16.49
Total gas mass $10^{9}{\rm\,M_{\odot}}$] 4.33
@@
Molecular gas mass $10^{9}{\rm\,M_{\odot}}$] 1.35
Neutral gas mass $10^{9}{\rm\,M_{\odot}}$] 1.06
Ionised gas mass $10^{9}{\rm\,M_{\odot}}$] 1.77
Ionised CGM gas mass $10^{9}{\rm\,M_{\odot}}$] 1.90
SFR $\rm\,M_{\odot}\,yr^{-1}$}] 43
@@ 4.1 Tidal tail
The first step of analysing the simulation is to visualise the merger process and the tidal tail, as it is one of the strongest signs of Haro 11 being a merger. In Figure 2 we present several projection maps of the gas and stars as part of a timeline of the simulation. These maps follow the formation of the tidal tail, from the first close interaction until they reach the current observed time of Haro 11, which is $\sim 10$ Myr into the second interaction; a time period of roughly $\sim 200$ Myr. Furthermore, the maps have Roman numerals annotated, corresponding to time markers in the figure of SFR as a function of time in Section 4.3, which highlights the burst of star formation during the close interactions.
This simulation matches the length and kinematics of the HI tail presented in Le Reste et al. (2024) using MeerKAT observations of the 21-cm line [^1] The reduced MeerKAT 21 cm data is publicly available at: https://archive-gw-1.kat.ac.za/public/repository/10.48479/7zn6-bw59/index.html. The tidal tail is extending in the same direction relative to the inner morphology as in Haro 11 (north-east) and has the same length $\sim 40$ kpc. Furthermore, we find that the relative HI mass inside the tidal tail is $\sim 36\,\%$, which is within the observed value $44^{+20}_{-19}$ %.
A velocity field map of the HI observed in emission only is presented in the left panel of Figure 3 (the equivalent plot for absorption can be found in Appendix A), and the HI velocity in our simulation is on the right. Most of the HI emission is red-shifted relative to the central galaxy, for which Le Reste et al. (2024) adopt the optical redshift from Bergvall et al. (2000); with similar systemic velocity as the ionised gas around the knots (see Menacho et al., 2021). Observations indicate the tidal tail is very slightly blue-shifted at a maximum speed of $\lesssim-20$$\rm{km\,s^{-1}}$}, while our simulation has a tidal tail with $\lesssim-50$$\rm{km\,s^{-1}}$}. This discrepancy is most directly explained by how the zero-velocity is set in the simulations, as there is no clear systemic velocity in a galaxy merger. Other possible factors are the viewing angle of the simulation or that the edges of the tidal tail are too faint to be detected by observations. We discuss this in more detail in Section   6.
@@
The stellar tidal tail is slightly shorter than the HI component ($\sim$30 kpc compared to $\sim$40 kpc). This is likely due to a difference in scale radius or truncation between the two components, i.e. the gas disc having a larger scale radius. We briefly explored the effects of truncation on the disc in Section 5.1. Additionally, the image of the stellar tidal tail shows clear shell structures around the galaxy. The simulation shows signs of these shells as well, which is encouraging for our model but is not directly applicable as a modelling constraint.
@@ 4.2 Inner morphology
In order to put our simulations into context, we first present an observation of the Haro 11 galaxy on the left in Figure 1. This was presented in Adamo et al. (2010) using HST in three wavebands filters; in particular they use F220W and F814W with the ACS/HRC instrument, and F435W with ACS/WFC. Additionally, the right panel of the figure shows a mock observation of our simulation using the post-processing radiative transfer code RASCAS (Michel-Dansac et al., 2020) to obtain the stellar continuum and apply the same HST filter profiles as the observations. Dust absorption follows the SMC law (e.g. Laursen et al., 2009) and assumes a dust-to-metal conversion factor, which increases with metallicity and cut-off at $T>10^{4}$ K. The key morphological features of Haro 11 are annotated: three distinct stellar knots (A, B, C) and a dusty arm going upwards from knot A to the right of knot B (the ’ear’). All of these morphological features can be identified in the mock observation as well, but knot A is not as readily apparent, which is discussed below.
The inner gas and stellar morphology, as well as temperature, of the galaxy at the current observed time is presented as projection maps in Figure 5. The high density gas around knot B and A correspond to areas of high stellar density. A disc arm with gas and stars is seen on the right side of the galaxy in each panel and is reminiscent of the ear observed in Haro 11. Compared to observations, the positions of these features are slightly mismatched in position relative to each other, but are broadly in the expected positions. Additionally, the dense gas that is overlapping with the stellar core on the left (knot C) is behind the core.
The main reason that knot A is not observed in our mock observations is due to the gaseous and dusty arm covering the stellar knot, but we can infer its presence from the stellar mass and young population. In Figure 6 we show a stellar map weighted by the inverse stellar age, as $M_{\rm*,\,young}=M_{*}\times(t/4\,{\rm Myr})^{-1.25}$, which is an approximate bolometric luminosity weighting; this highlights the presence of young, luminous stars. This figure shows that there is significant mass of young stars around knot A, which would be visible in the absence of extinction. The reason why knot A is more obscured, and other reason why knot A is abscent in our mock image, is discussed in more detail in Section 6. Additionally, in Figure 8 we present a projection map of the median stellar ages, from which it is also evident that all of the knots, including A, contain many young stars. From this, we infer that knot A likely reside within the disc arm (which also makes up the ear) of the prograde progenitor galaxy, which has its centre at knot B. The knots’ properties are presented in more detail in Section 4.4.
@@ 4.3 Star formation rate
Observations show that Haro 11 is currently experiencing a burst of star formation, which is attributed to an undergoing merger. We present the SFR as a function of time in Figure 7, annotated with Roman numerals as timestamps to the timeline images in Figure 2 and with shaded areas for the time of pericentre passages. The figure shows two peaks in the SFR that are associated with close interaction between the two galaxies. The SFR burst is thus likely caused by interaction effects, e.g. gas compression due to tidal forces, cloud-on-cloud collision, tidal torques that drive gas towards the centres of the galaxies (e.g. Renaud et al., 2014b, 2019).
The current day SFR of our fiducial simulation match with the observed values, ${\rm SFR}\lesssim 30{\rm\,M_{\odot}\,yr^{-1}}$ (see Section 2 for references). Furthermore, the starburst periods in our simulations can be readily compared to observations of stellar populations with distinct ages, which helps constrain the star formation history. The peaks indicate that these stellar populations should have ages of roughly 200250200-250 Myr, from the first interaction, and a few Myr, from the current interaction. We will explore this further in the next section by comparing stellar analysis models of the populations in the three knots.
Figure 7: The star formation rate of the simulation. The time axis is set that the current observed time is at 0 Myr. Only the last $\sim 500$ Myr are shown, as there is no interaction between galaxies before this and the galaxies have reached a roughly constant SFR (see Section 3.4 for details). Specific time periods are marked with Roman numerals, corresponding to the projection maps with same number in Figure 2. The grey shaded areas are the estimated pericentre periods of interaction. The starbursts at 0 Myr and 250 Myr correspond to close, gravitational interactions. The diamond marks the time at which the galaxy best resembles Haro 11, following the criteria in Section 3.2, and the square represents the SFR range mentioned in Section 2. The galaxy is in the middle of its second starburst event, with a ${\rm SFR}=43{\rm\,M_{\odot}\,yr^{-1}}$.
@@ 4.4 Knot properties
The distribution of mass and median ages of the clusters, divided into their knot affiliation, are presented in Figure 9. As mentioned in Section 4.3, we notice two massive populations of star clusters that form during the starburst peaks that coincide with the close interactions. These groups of population ages, $\sim 10^{6}$ Myr and $\sim 10^{8}$ Myr, have been found in stellar population synthesis models of Haro 11 (Chandar et al., 2023; Papaderos & Östlin, 2023). We recover in the simulation star clusters as massive as those found in the observations at the youngest age, while the majority of the old population would remain undetected at masses below a few times $10^{5}{\rm\,M_{\odot}}$; as can be seen from the 90% completeness limit from Sirressi et al. (2022), presented as a dark line. The most massive clusters found in our simulations $\sim 10^{8}{\rm\,M_{\odot}}$ are not found in the observations, but could stem from insufficient mass resolution, the cluster identification, or other, more spurious, effects. For future work, Ejdetjärn et al. (in prep), we plan to perform more extensive analysis of the cluster formation within our high-resolution simulations of the same setup.
@@ 4.5 Kinematics
In Ö15, the authors made a comparison between the position-velocity (PV) diagram of the Antennae galaxies and Haro 11, which they noted had a similar shape and velocity magnitude. They acquired data using the CIGALE instrument and archived data for the Antennae. Their data covered a straight line going from knot C to A, and extending a few kpc beyond them. We follow their method and present our own PV diagram in Figure 11. The $x$-axis is determined as the projected distance along the dark line on the left plot in Figure 10. The left/right plot in the figure shows the simulation/observational data, which have very similar shapes with peaks and valleys at roughly the same distance intervals. We highlight these similarities, with a broad grey line in both plots, which is the mean velocity of the simulation at each distance in bins of 0.5 kpc. This figure also highlights that several of the simulations cells attain much higher speeds, as much as twice as fast, as the observations. However, the mean velocity (grey line) follows the same sine-like shape and is within the limits of the observational data.
@@ 5.1 Producing one tidal tail
Additionally, truncating the gas disc also has an effect on the resulting morphology of the tidal tails. In the centre row of Figure 12 we show an initial galaxy with retrograde and prograde motions (left), a thicker disc for the retrograde (middle), and a truncated disc for the retrograde (right). The thicker disc can be seen to yield a slightly weaker second tail and have more tidal features surrounding it. The truncated disc also shows a more diffuse second tail and possibly some small changes to the main tail.
From our simulation tests, we learned that the length of the tidal tail puts a constraint on the time since the first interaction, when the tail first forms. In order for the gas tail to reach the observed 40 kpc, it would need roughly 200 Myr to expand. For comparison, a shorter interaction time of $\sim 100$ Myr would yield a length of $\sim$20 kpc. This time might be a lower limit, as the observed expansion speed of tidal tail is slightly lower in the HI observations. Furthermore, the presence of a tidal tail this long also substantiates that Haro 11 is at least in its second close interaction.